OpenStudy (swaqqout_kid1):

Solve for x. x = -8 x = 56 x = -10 x = -6 @ttb123456789

OpenStudy (anonymous):

What is the equation? @SwaqqOut_KID1

OpenStudy (swaqqout_kid1):

https://media.studyisland.com/cgi-bin/mimetex.cgi? \frac{x}{4}\%20+\%208\%20=\%206

OpenStudy (anonymous):

equation please i will help

OpenStudy (anonymous):

where is it?

OpenStudy (emmamink):

Where the heck is the thing?

OpenStudy (ageta):

Sorry, I don't understand your input.

OpenStudy (anonymous):

draw it

OpenStudy (swaqqout_kid1):

how

OpenStudy (anonymous):

click draw

OpenStudy (swaqqout_kid1):



OpenStudy (anonymous):

x/4+8=6 x/4=6-8 x=(6-8)*4

OpenStudy (anonymous):

https://media.studyisland.com/cgi-bin/mimetex.cgi? \frac{x}{4}\%20+\%208\%20=\%206

OpenStudy (anonymous):

Copy the whole thing, and paste it.

OpenStudy (anonymous):

siplify

OpenStudy (anonymous):

write what you will get

OpenStudy (anonymous):

medal

OpenStudy (swaqqout_kid1):

-8

OpenStudy (anonymous):

let me check -2*4=-8 YES